Katy:

Find the area of the shaded region. Leave your answer in terms of π. (4+4π) ft^2 (4-π) ft^2 (4+π) ft^2 (4-4π) ft^2

justjm:

It is asking you to find the area between the square and the inscribed circle. Therefore, you must find \(A_{shaded~region}=A_{square}-A_{circle}\) •To find the area of the square, it is already giving you the side lengths. A=s^2 •To find the area of the inscribed circle, you may need to find the radius, as it is not explicitly given. Since the circle is inscribed in the square, the diameter of the circle is equivalent to the side length of the square. Hence, the radius would be half the side length of the square, so r=1. Now you can find A=πr^2

justjm:

I cannot give away direct answers but I can lead you to the correct choice...can you tell me the area of the square and the area of the circle?

justjm:

The area of a square is s^2 where s is side. Here the side length is 2, so the area is _______ The area of a circle is πr^2, here r=1 because the diameter is equal to the sidelength, which is 2. So the area of the inscribed circle is _______

justjm:

Yes it is one of those two

justjm:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Katy (4-π) ft^2 (4-4π) ft^2 Which one? \(\color{#0cbb34}{\text{End of Quote}}\) I can't just give it away, which one do you think?

justjm:

No, your area of the circle is wrong, What made you get 4 pi?

InsatiableSuffering:

Remember the area of a circle is A=pi*r^2. r is radius. In order to find radius, you have to take half of the diameter. You know what the diameter is. Take half of that, square the radius, and mulitply by pi.

justjm:

Yes